# Enhancing Peer Review in Astronomy: A Machine Learning and Optimization Approach to Reviewer Assignments for ALMA

### v1

@@ Abstract
The increasing volume of papers and proposals undergoing peer review emphasizes the pressing need for greater automation to effectively manage the growing scale. In this study, we present the deployment and evaluation of machine learning and optimization techniques for assigning proposals to reviewers that was developed for the Atacama Large Millimeter/submillimeter Array (ALMA) during the Cycle 10 Call for Proposals issued in 2023. By utilizing topic modeling algorithms, we identify the proposal topics and assess reviewers’ expertise based on their historical ALMA proposal submissions. We then apply an adapted version of the assignment optimization algorithm from PeerReview4All (Stelmakh et al., 2021a) to maximize the alignment between proposal topics and reviewer expertise. Our evaluation shows a significant improvement in matching reviewer expertise: the median similarity score between the proposal topic and reviewer expertise increased by 51 percentage points compared to the previous cycle, and the percentage of reviewers reporting expertise in their assigned proposals rose by 20 percentage points. Furthermore, the assignment process proved highly effective in that no proposals required reassignment due to significant mismatches, resulting in a savings of 3 to 5 days of manual effort.
software: gensim (Řehůřek & Sojka, 2010), gurobi (Gurobi Optimization, LLC, 2023), matplotlib (Hunter, 2007), PeerReview4All (Stelmakh et al., 2021a), spaCy (Honnibal et al., 2020)
@@ 1 Introduction
Peer review is the cornerstone of modern scientific research, playing a crucial role in the allocation of billions in grant funding and providing access to high-cost facilities such as supercomputers and telescopes, both ground- and space-based. Traditionally, this process relies on an invited panel of experts to evaluate and recommend which proposals should be accepted. This model has been widely adopted across the scientific community, including by major astronomical observatories. One prominent example is the Atacama Large Millimeter/submillimeter Array (ALMA), which utilized panel-based peer review when it began science operations in 2011. As the largest ground-based observatory and the most sensitive telescope ever constructed for high-resolution imaging of the submillimeter sky, ALMA is in high demand, making an efficient peer review process essential to its operations.
The increasing volume of proposals submitted to major observatories has put the traditional panel-based peer review process under strain. The Hubble Space Telescope (HST), the James Web Space Telescope (JWST), the European Southern Observatory (ESO), and ALMA for example receive between 1000 to 2000 proposals per year. Managing this large volume requires an ever-growing number of reviewers, complicating the logistics of organizing panel meetings, which can require 200 or more participants. These challenges have prompted some observatories to evolve the peer review process from the traditional panel approach to allow for proposal review without a face-to-face or virtual meeting of the reviewers. HST and JWST now use external reviewers in this manner for the smaller time requests, and distributed peer review (Merrifield & Saari, 2009) is in use at Gemini for Fast Turnaround proposals (Andersen et al., 2019), the European Southern Observatory (Patat et al., 2019; Jerabkova et al., 2023), and ALMA (Donovan Meyer et al., 2022), where each proposal team designates one person to participate in the review process.
In addition to logistical challenges, the assignment of proposals to reviewers has also become increasingly complex due to the large volume of submissions. In smaller systems, this task can often be handled manually, but for the largest observatories, automated approaches are necessary. Commonly, proposals and reviewers are tagged with topical keywords to facilitate matching based on shared expertise. However, this method has limitations, particularly in the accurate assignment of keywords to both proposals and reviewers (Strolger et al., 2023). With the increasing use of distributed peer review processes, emphasis is also placed on matching the proposals to the reviewers expertise since the reviewers will not have a group discussion of the proposals. As a result, HST, JWST, and ESO have begun incorporating machine learning and optimization techniques to improve the matching of proposals to reviewer’s expertise (Strolger et al., 2017, 2023; Kerzendorf et al., 2020).
These advanced algorithms typically work in two phases (see Section 3 in Shah, 2022). First, machine learning methods are used to quantify the similarity between the subject matter of a proposal and the expertise of the reviewer. Based on these similarities, proposals are then assigned to reviewers through optimization of a pre-determined metric. Notably, automated machine learning algorithms have been integrated into the assignment processes of HST and JWST (Strolger et al., 2017, 2023) and in a pilot initiative at ESO (Patat et al., 2019; Kerzendorf et al., 2020). Following these examples, ALMA introduced machine learning and optimization techniques into its review process in 2023 (ALMA Cycle 10).
This paper details ALMA’s deployment of this new assignment process and evaluates its performance. Our main contributions are as follows:
  • Deployment. We deploy machine learning and optimization techniques to assign reviewers to the submitted proposals.
    • We use topic modeling to compute a predicted expertise (termed “similarity,” a number between 0 and 1) of each reviewer for each submitted proposal.
    • We then employ the PeerReview4All optimization algorithm (Stelmakh et al., 2021a) to optimally assign reviewers to the proposals, taking the aforementioned predicted expertise as input.
  • Evaluation. We evaluate the quality of the deployed assignment in several ways.
    • In the current cycle, assigned reviewer-paper pairs have a median similarity score of 0.71, up from 0.20 in the previous cycle.
    • We also conducted a reviewer survey, where we find that reviewers indicated a higher level of expertise in their proposal assignments with the new algorithm. The percentage of reviewers claiming expertise in their assigned proposals rose from 45% in the previous cycle to 65% in this cycle, while those indicating little or no knowledge decreased from 10% to 5%.
    • The process was highly effective, as no manual adjustments were required to address any significant mismatches between reviewer expertise and proposal topics, as compared to 3 to 5 days of manual effort previously.
The paper is organized as follows. Section 2 provides an overview of the review process used by ALMA in previous cycles and the motivation to implement a new approach. In Section 3, we describe related work in astronomy and other fields that have implemented machine learning and optimization techniques that was incorporated into ALMA’s efforts. Section 4 describes the deployment of the new system, including the use of machine learning to infer the topic of individual proposals and the reviewer’s expertise, and the assignment of proposals to individual reviewers. Section 5 evaluates the results of the new process, including back testing the results to previous cycles and presenting the results from actual use in Cycle 10. Section 6 discusses the results and summarizes the conclusions.
2 Overview of ALMA Proposal Assignment in Previous Cycles
ALMA began offering observing time in 2011 and solicits proposals annually. The number of submitted proposals increased steadily from 919 proposals in 2011 (Cycle 0) to a peak of 1836 in 2018 (Cycle 6), and currently about 1700 proposals are submitted annually. PIs submit their proposals under one of five scientific categories. [^1] Category 1: Cosmology and the high redshift universe. Category 2: Galaxies and galactic nuclei. Category 3: Interstellar medium, star formation and astrochemistry. Category 4: Circumstellar disks, exoplanets and the solar system. Category 5: Stellar evolution and the Sun. Within each category, they also select one or two keywords from a list provided by ALMA that best describe the topic of their proposal.
Initially, ALMA employed a panel-based review process, where panels of 6 to 9 experts for each of the five scientific categories evaluated around 100 proposals each. Multiple panels were formed per category as needed depending on the number of submitted proposals. However, this system became strained as the number of proposals increased, prompting ALMA to implement a distributed peer review process. This system was first used in 2019 (Cycle 7) for a “supplemental” call to fill time on the Morita Array (also known as the Atacama Compact Array) that was not allocated in the main call (Carpenter et al., 2020). ALMA used distributed peer review in the main call for the first time in 2021 (Cycle 8) for a subset of proposals, and adopted the distributed process on all proposals except for ALMA Large Programs in 2022 (Cycle 9).
In the distributed peer review system, the Principal Investigator (PI) of each proposal nominates a team member to serve as a reviewer. Each reviewer assesses 10 proposals, and each proposal receives 10 reviews. Reviewers rank the proposals from 1 (strongest) to 10 (weakest), and the final overall rankings are based on the averaged individual rankings. The process is detailed by Donovan Meyer et al. (2022). The number of proposals evaluated by this process was 1497, 1729, and 1635 in Cycles 8, 9, and 10, respectively.
@@ 2 Overview of ALMA Proposal Assignment in Previous Cycles
Initially, ALMA used a “category” and “keyword” approach to match proposals with reviewers in the distributed system. Reviewers specified their scientific expertise using the same list of keywords used by PIs to characterize their proposals. A series of “rules” were specified that established the priority order on how to assign proposals to reviewers (see Donovan Meyer et al., 2022). The top priority was to assign proposals where one or more of the proposal keywords specified by the PI overlapped with the expertise keywords specified by the reviewer. Proposals that had the fewest available expert reviewers were assigned first. The assignment process successfully aligned the keywords of the assigned proposals to the reviewer’s expertise keywords in about 95% of cases. To gauge the success of the assignment process from the reviewer’s perspective, each reviewer had the option to rate their expertise on the proposal, with possible options of “My field of expertise”, “Some knowledge”, or “Little or no knowledge”. Reviewers indicated that had at least some knowledge in 90% of the assigned proposals and little or no knowledge in just 10% of the assignments (Donovan Meyer et al., 2022).
@@
Despite the success of the category and keyword approach, some limitations persist. In the 10% of cases where reviewers reported little or no expertise, the proposal keywords often matched the reviewer’s expertise keywords (Donovan Meyer et al., 2022). This discrepancy could arise from several factors: the keywords may be too broad, reviewers might have selected incorrect keywords, they may be conservative in assessing their expertise, or they might be relatively inexperienced, as ALMA allows student reviewers. Additionally, keywords selected by reviewers might not fully capture their expertise, making it difficult to identify all suitable reviewers for a given proposal. Another challenge is the logistical need to assign 10 proposals to each reviewer and ensure each proposal is reviewed 10 times. The assignment algorithm had some deficiencies in that not all reviewers received proposals aligned with their expertise, and in extreme cases, some reviewers were assigned proposals outside their field (e.g., a circumstellar disk proposal assigned to an extragalactic expert). These mismatches were resolved by identifying reviewers to swap individual assignments. However, this process took 3 to 5 days of manual effort to complete, and was the most time consuming task in the assignment process.
@@
ALMA set out to implement a new approach to assign proposals to reviewers with three objectives. First, the proposal assignment process should result in an overall increase in the expertise as measured by the similarity between the proposal topic and the reviewer’s expertise. Second, reviewers should indicate a higher level of expertise in their assignments. And finally, the process needs to be automated and robust to reduce the amount of time spent on manual re-assignments.
3 Related work
Other observatories and scientific fields face similar challenges as ALMA in that they receive hundreds if not thousands of proposals or papers that needed to be reviewed. Automated techniques are needed to assign proposals or papers to reviewers in an optimal manner. This section describes some of these previous efforts. While not an exhaustive review of the subject, we highlight previous efforts that most impacted ALMA’s implementation of machine learning and assignment optimization. We begin by discussing two prior deployments for proposal reviewing in astronomy, followed by deployments for paper review in computer science, and conclude the section by noting differences with the present work.
3.1 Space Telescope Science Institute
The Space Telescope Science Institute has developed a tool called the Proposal Auto-Categorizer and Manager for Time Allocation Review (PACMan; Strolger et al., 2017, 2023) that provides a number of services for proposal evaluation for the Hubble Space Telescope (HST) and the James Web Space Telescope (JWST). PACMan functionality includes using machine learning techniques to classify proposals into broad topics, inferring the expertise of reviewers based on their abstracts from their papers in the Astrophysics Data System (ADS), [^2] https://ui.adsabs.harvard.edu matching proposals to reviewers, and identifying conflicts of interest, among other tasks. Proposal classification and reviewer’s expertise are determined using the Term Frequency-Inverse Document Frequency (TF-IDF) model. Their results indicate that the top topical category inferred by PACMan agrees with the manual classification in 84% of the cases on average across categories. Similarly, the category of scientific expertise inferred for the reviewers has a high correspondence with the expertise provided by the reviewers.
3.2 European Southern Observatory
The European Southern Observatory (ESO) implemented machine learning techniques (Kerzendorf, 2019; Kerzendorf et al., 2020) for their pilot run of distributed peer review (Patat et al., 2019, see also Jerabkova et al. 2023), which was run in parallel with a traditional panel review. The topic of each proposal was inferred from the proposal text using the TF-IDF model, while the reviewer expertise was established based on their publication history in ADS. The dot product of the proposal vector and the reviewer expertise (“cosine distance”) then provided a measurement of the expertise. The proposals were then assigned to reviewers by optimizing the similarities of the assigned proposals. The results show a high correlation between the predicted expertise and the self-reported expertise from the reviewers (Kerzendorf et al., 2020).
3.3 Computer Science
Computer science has used distributed peer review for well over a decade to select papers submitted to conferences (for an overview of the review process, see Shah, 2022). The largest conferences can have over 13,000 submissions (NeurIPS, 2023), and therefore automated tools are essential to handle so many papers. These conferences review full papers (and not just abstracts), are frequently a terminal venue for publication, and are rated at par or better than journals.
Here is a short summary of the review process in such conferences. After papers are submitted to the conference, they are assigned to reviewers in a (semi-)automated fashion in a manner similar to that adopted at ALMA. First, a similarity score is computed via one or more of the following methods: automated text-matching (Charlin et al., 2012; Mimno & McCallum, 2007; Cohan et al., 2020), computing the intersection between reviewer- and author-chosen subject areas, and by allowing reviewers to manually specify their interest and/or expertise for various papers (called “bidding”; see Fiez et al., 2020). If more than one of these methods are used, then the outcomes are combined to obtain a single similarity score for each reviewer-paper pair. Subsequently, an optimization procedure assigns reviewers to papers. The metric that is optimized depends on the conference, and includes maximizing the total sum of similarity scores of the assignments (Goldsmith & Sloan, 2007; Charlin et al., 2012), or maximizing the similarities for the papers that are worst-off in terms of available expertise (Stelmakh et al., 2021a; Kobren et al., 2019). We refer the reader to Stelmakh et al. (2023), Saveski et al. (2023), and Stelmakh et al. (2021a) for some evaluations of the reviewer assignment procedures in computer science.
After initial reviews are submitted by the reviewers, authors get a chance to rebut any incorrect statements in the reviews (Gao et al., 2019; Liu et al., 2023). The reviewers get to read the rebuttal, optionally update their reviews, and engage in a discussion with each other on a typed forum (Stelmakh et al., 2020; Rastogi et al., 2024). A meta-reviewer (equivalent of an associate editor in journals) is also involved in the discussion who eventually makes a recommendation to the program chair on which papers should be accepted. Typically, about 20%–25% of the submitted papers are accepted to the conference.
3.4 Influence on ALMA
HST, JWST, ESO, and computer science have shown the viability of machine learning and optimization techniques that inspired ALMA’s efforts. ALMA’s implementation differs from those used by other observatories in a few key ways. First, the reviewer’s expertise was established based on their ALMA proposal history rather than ADS data. Each reviewer is guaranteed to be on at least one ALMA proposal by design of the system, and a reviewer’s proposals can be identified unambiguously. Second, proposal categories and keywords are combined with reviewers’ expertise keywords to refine assignment options, reducing the number of spurious matches that machine learning alone might generate. Lastly, we adopted a computer science optimization algorithm for assigning proposals, with a few adjustments to fit ALMA’s distributed peer review model. Section 4 outlines the deployment of this new assignment process in detail.
4 Deployment
@@ 4.1 Topic Modeling
The first step in the proposal assignment process is to determine the topics of each proposal using natural language processing. The PDF files for submitted ALMA proposals from Cycle 1 onwards were converted to ASCII text using the python library pdftotext. [^3] https://github.com/jalan/pdftotext The technical justification portion of the proposal that describes and justifies the setup of the observations is removed given it is largely technical in nature and much of the text will appear similar between proposals. The ASCII text for the scientific justification was then combined with the proposal title and proposal abstract stored in the ALMA databases to define the proposal text. Pre-processing steps were applied to the resulting files using spaCy (Honnibal et al., 2020) and custom functions to define and remove stop words, standardize commonly used acronyms, convert British spelling of words to American spelling, and lemmitization. This process created a “bag of words” for each proposal.
Many methods are available to classify documents from the bag of words. We mainly explored the use of TF-IDF, Latent Semantic Indexing, and Latent Dirichlet Allocation (LDA). All three methods gave satisfactory and similar results based on how well they can find similar documents which are known apriori to share the same scientific category and keywords. We ultimately settled on the LDA method (Blei et al., 2003) since we found for our application that it was able to match proposals to within the same scientific category at a slightly higher rate than TF-IDF and LSA.
@@
Here, $S(\vec{p_{i}},\vec{p_{j}})$ ranges between 0 and 1, where a value of 0 indicates that the topics of the two proposals are dissimilar and a value of 1 indicates the topics are the same. For each proposal, the 10 proposals that have the highest similarity were identified, and we computed the fraction of those proposals that were submitted to the same scientific category as selected by the PI. While this is by no means a perfect measure of the accuracy of the algorithm since there are ambiguities in the definitions of the scientific categories, it does provide a first order indication of how well the algorithm performs since proposals in the same scientific category are expected to have similar content.
Figure 2 shows the results for the Cycle 10 proposals. As an example, for proposals submitted to Category 1, 80% of the matched proposals were also in Category 1, 17% in Category 2, 2% in Category 3, 0% in Category 4, and 1% in Category 5. The high overlap between Categories 1 and 2 is expected given both contain galaxy proposals and there is no precise division by redshift. More relevant is that only 3% of the proposals were matched in Categories 3, 4, and 5 combined. Even the 1% matching in Category 5 is not necessarily spurious, since this category contains transient proposals which overlaps with Gamma Ray Bursts in Category 1. Category 5 is the one category that has significant overlap with all other categories. This reflects that large range of topics in this category (e.g., the Sun, transients, pulsars, evolved stars, blackholes), and it also contains the fewest overall proposals, so there may not be 10 proposals that are closely matched. The main conclusion is that while most of closest matched proposals are in the same category as the submitted proposal, there are potential false matches.

### v2

@@ Abstract
The increasing volume of papers and proposals that undergo peer review emphasizes the pressing need for greater automation to effectively manage the growing scale. In this study, we present the deployment and evaluation of machine learning and optimization techniques to assign proposals to reviewers that were developed for the Atacama Large Millimeter/submillimeter Array (ALMA) during the Cycle 10 Call for Proposals issued in 2023. Using topic modeling algorithms, we identify the proposal topics and assess reviewers’ expertise based on their previous ALMA proposal submissions. We then apply an adapted version of the assignment optimization algorithm from PeerReview4All (Stelmakh et al., 2021a) to maximize the alignment between proposal topics and reviewer expertise. Our evaluation shows a significant improvement in matching reviewer expertise: the median similarity score between the proposal topic and reviewer expertise increased by 51 percentage points compared to the previous cycle, and the percentage of reviewers reporting expertise in their assigned proposals rose by 20 percentage points. Furthermore, the assignment process proved highly effective in that no proposals required reassignment due to significant mismatches, resulting in a savings of 3 to 5 days of manual effort.
software: gensim (Řehůřek & Sojka, 2010), gurobi (Gurobi Optimization, LLC, 2023), PeerReview4All (Stelmakh et al., 2021a), spaCy (Honnibal et al., 2020)
@@ 1 Introduction
Peer review is the cornerstone of modern scientific research, playing a crucial role in the allocation of billions in grant funding and providing access to high-cost facilities such as supercomputers and telescopes, both ground- and space-based. Traditionally, this process relies on an invited panel of experts to evaluate and recommend which proposals should be accepted. This model has been widely adopted across the scientific community, including most major astronomical observatories. A prominent example is the Atacama Large Millimeter/submillimeter Array (ALMA), which implemented panel-based peer review when it began science operations in 2011. As the largest ground-based observatory and the most sensitive telescope ever constructed for high-resolution imaging of the submillimeter sky, ALMA’s high demand necessitates an efficient peer review process.
The number of proposals submitted to ALMA has steadily increased, from 919 in 2011 (Cycle 0) to a peak of 1836 in 2018 (Cycle 6). Currently, approximately 1700 proposals are submitted annually. This growing volume created logistical challenges for the traditional panel-based process, prompting ALMA to adopt a distributed peer review system. In this system, each proposal team designates one member to review ten proposals. ALMA first implemented distributed peer review in 2019 (Cycle 7) for a ”supplemental” call to fill unallocated time on the Morita Array (Carpenter et al., 2020). Distributed peer review was then used for a subset of proposals in the main call in 2021 (Cycle 8) and fully adopted for all proposals, except Large Programs, in 2022 (Cycle 9). A comprehensive description of ALMA’s distributed peer review process is provided by Donovan Meyer et al. (2022).
@@
While distributed peer review addressed logistics challenges, it introduced complexities in assigning proposals to reviewers. Initially, assignments were based on common categories and keywords between the submitted proposals and the expertise self-identified by the reviewers. The top priority was to assign proposals where one or more of the proposal keywords specified by the principal investigator (PI) overlapped with the expertise keywords specified by the reviewer. Proposals that had the fewest available expert reviewers were assigned first. This process successfully aligned proposal and reviewer keywords in about 95% of cases. To gauge the success of the assignment process from the reviewer’s perspective, each reviewer had the option to rate their expertise on their assignments, with possible options of “My field of expertise”, “Some knowledge”, or “Little or no knowledge”. Reviewers reported at least some knowledge in 90% of the assigned proposals, with only 10% categorized as “little or no knowledge” (Donovan Meyer et al., 2022).
@@
Despite its successes, the keyword-based approach faced limitations. In the 10% of cases where reviewers reported little or no expertise, the assigned proposal keywords often still matched the reviewer’s expertise (Donovan Meyer et al., 2022). This discrepancy could result from broad or imprecise keywords, conservative self-assessments of expertise by the reviewers, or inexperience among reviewers, particularly student participants. Additionally, the keywords provided by reviewers may not fully capture their expertise, complicating the identification of suitable matches. Another challenge is the logistical need to assign 10 proposals to each reviewer and ensure each proposal is reviewed 10 times. In some instances, reviewers received proposals outside their primary field (e.g., a circumstellar disk proposal assigned to an extragalactic expert). Resolving these mismatches required three to five days of manual effort to swap assignments, making it the most time-consuming step in the assignment process.
Other observatories and scientific disciplines have experienced similar challenges and have turned to advanced machine learning and optimization techniques to enhance the efficiency of proposal assignments. These methods typically operate in two phases (see Section 3 in Shah, 2022): (1) machine learning methods quantify the similarity between a proposal’s subject matter and a reviewer’s expertise, and (2) optimization algorithms assign proposals to reviewers based on a defined metric.
For example, the Space Telescope Science Institute (STScI) employs machine learning techniques to classify Hubble Space Telescope (HST) and James Webb Space Telescope (JWST) proposals into broad topics. Reviewer expertise is inferred from their publication abstracts in the Astrophysics Data System (ADS), [^1] https://ui.adsabs.harvard.edu enabling effective reviewer-proposal matching (Strolger et al., 2017, 2023). Similarly, the European Southern Observatory (ESO) implemented machine learning techniques for distributed peer review (Kerzendorf, 2019; Kerzendorf et al., 2020; Jerabkova et al., 2023). Proposal topics are inferred from proposal text, while reviewer expertise is determined from publication histories in ADS.
In computer science, machine learning and optimization techniques have long been integral to assigning papers submitted to conferences during distributed peer review. Major conferences, such as NeurIPS, handle over 13,000 submissions (NeurIPS, 2023), necessitating automated tools. These tools compute similarity scores using methods such as automated text matching (Charlin et al., 2012; Mimno & McCallum, 2007; Cohan et al., 2020), intersection of reviewer- and author-selected subject areas, and manual reviewer “bidding” to express preferences (Fiez et al., 2020). Optimization algorithms then assign reviewers to papers, maximizing similarity scores or addressing fairness concerns (Stelmakh et al., 2021a; Kobren et al., 2019; Stelmakh et al., 2023).
@@
Inspired by the efforts at other observatories and in computer science, ALMA developed a new proposal assignment approach with three primary objectives. First, the proposal assignment process should result in an overall increase in the expertise as measured by the similarity between the proposal topic and the reviewer’s expertise. Second, reviewers should indicate a higher level of expertise in their assignments. Finally, the process needs to be automated and robust to reduce the amount of time spent on manual reassignments.
This paper presents the implementation and evaluation of the new assignment system. Section 2 describes the methodology for computing similarities between a reviewer’s expertise and the submitted proposals, as well as optimizing assignments using an adapted version of the PeerReview4All algorithm (Stelmakh et al., 2021a). Section 3 establishes the utility of the similarity measures by correlating Cycle 8 and Cycle 9 assignments with reviewers’ self-reported expertise and evaluates the performance of the overall assignment process following its full deployment in Cycle 10. Section 4 provides a summary of our findings and suggestions for future improvements.
2 Implementation
@@ 2.1 Topic Modeling
The first step in the proposal assignment process is to determine the topics of each proposal using natural language processing. The PDF files for submitted ALMA proposals from Cycle 1 onwards were converted to ASCII text using the python library pdftotext. [^2] https://github.com/jalan/pdftotext The technical justification portion of the proposal that describes and justifies the setup of the observations is removed given it is largely technical in nature and much of the text will appear similar between proposals. The ASCII text for the scientific justification was then combined with the proposal title and proposal abstract stored in the ALMA databases to define the proposal text. Pre-processing steps were applied to the resulting files using spaCy (Honnibal et al., 2020) and custom functions to define and remove stop words, standardize commonly used acronyms, convert British spelling of words to American spelling, and lemmatization. This process created a “bag of words” for each proposal.
Many methods are available to classify documents from the bag of words. We explored the use of Term Frequency-Inverse Document Frequency (TF-IDF), Latent Semantic Indexing (LSI), and Latent Dirichlet Allocation (LDA). All three methods gave satisfactory and similar results based on how well they can find similar documents which are known apriori to share the same scientific category and keywords. We ultimately settled on the LDA method (Blei et al., 2003) since we found for our application that it was able to match proposals to within the same scientific category at a slightly higher rate than TF-IDF and LSI.
@@
Here, $S(\vec{p_{i}},\vec{p_{j}})$ measures the cosine similarity between the topic vectors $i$ and $j$, ranging from 0 (complete dissimilar) to 1 (identical topics). However, the relative values of the similarities are more meaningful than the absolute values. For each proposal, the 10 proposals that have the highest similarity were identified, and we computed the fraction of those proposals that were submitted to the same scientific category as selected by the PI. While this is by no means a perfect measure of the accuracy of the algorithm since there are ambiguities in the definitions of the scientific categories, it does provide a first order indication of how well the algorithm performs since proposals in the same scientific category are expected to have similar content.
Figure 2 shows the results for the Cycle 10 proposals. As an example, for proposals submitted to Category 1, 80% of the matched proposals were also in Category 1, 17% in Category 2, 2% in Category 3, 0% in Category 4, and 1% in Category 5. The high overlap between Categories 1 and 2 is expected given both contain galaxy proposals and there is no precise division by redshift. More relevant is that only 3% of the proposals were matched in Categories 3, 4, and 5 combined. Even the 1% matching in Category 5 is not necessarily spurious, since this category contains transient proposals which overlap with Gamma Ray Bursts in Category 1. Category 5 is the one category that has significant overlap with all other categories. This reflects that large range of topics in this category (e.g., the Sun, transients, pulsars, evolved stars, blackholes), and it also contains the fewest overall proposals, so there may not be 10 proposals that are closely matched. The main conclusion is that while most of closest matched proposals are in the same category as the submitted proposal, there are potential false matches.